Banyan Co.’s common stock currently sells for $52.75 per share. The growth rate is a constant 10.5%, and the company has an expected dividend yield of 5%. The expected long-run dividend payout ratio is 25%, and the expected return on equity (ROE) is 14%. New stock can be sold to the public at the current price, but a flotation cost of 15% would be incurred. What would be the cost of new equity? Round your answer to two decimal places. Do not round your intermediate calculations.
Answer #1

Price of share after flotation costs =52.75*(1-15%) =44.8375

Expected dividend =5%*44.8375 =2.241875

Cost of new equity =Expected dividend/Current price +growth rate =2.241875/44.8375+10.5%= 15.50%